OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4. 81x4 12x5 81x5 12x4

OpenStudy (anonymous):

\[3^4.x^4\] Solve it or not???

ganeshie8 (ganeshie8):

\[\huge{(ab)^n = a^n.b^n}\]

OpenStudy (anonymous):

\[81{4}\]

ganeshie8 (ganeshie8):

latex error is it ? its \[81x^4\]

OpenStudy (goformit100):

81x4

OpenStudy (anonymous):

its 81 to the power of 4

OpenStudy (anonymous):

Then you are totally wrong @rajfk35 ...

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

\[3^4.x^4 = 81.x^4\]